OpenStudy (anonymous):

The area of this square is 17 inches squared. Choose the correct length of 1 side. A. Square Root Of 4.25 in. B. Square Root Of 8.5 in. C. Square Root Of 34 in. D. Square Root Of 17 in.

OpenStudy (anonymous):

 This is the square

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

just in case we didnt know what a square was?

OpenStudy (anonymous):

haha good one

OpenStudy (zzr0ck3r):

ok you have the area = 17 and area is given by \(A=x^2\) where x is the length of one side so we have \(17=x^2\) which implies \(x=\pm\sqrt{17}\) but since its a length we only take the positive root thus length \(=x=\sqrt{17}\)

OpenStudy (anonymous):

THANX!!! @zzr0ck3r

OpenStudy (zzr0ck3r):

npz